answersLogoWhite

0

The probability of rolling one 6 is ( 1/6 ).

The probability of rolling seven 6's is ( 1/6 )7 = 3.572 x 10-6 = 0.0003572 percent

User Avatar

Wiki User

∙ 17y ago

What else can I help you with?

Related Questions

What are the chances of rolling doubles 7 times in a row?

The probability of rolling doubles on a fair six-sided die is 1/6. To roll doubles 7 times in a row, you would need to multiply this probability by itself 7 times, resulting in (1/6)^7. This equals approximately 1 in 78,364,164,096, which means the chances of rolling doubles 7 times in a row are extremely low.


What are the chances of rolling a six on a die six times in a row?

The probability of rolling a six on a die 6 times in a row is 1 in 46,656, or 0.000021433. The calculate that, the probability of rolling one 6 is 1 in 6. Simply raise that to the sixth power to get 1 in 46,656.


What are the odds of rolling a 3 4 times in a row?

The probability of rolling a 3 on a die are 1/6. Since each roll is an independent event, the probability of rolling a 3 4 times in a row is (1/6)4 or 1/1296.The odds of something happening are the chances of it happening vs. the chances of it not happening. The chance of not rolling a 3 4 times in a row is 1-(1/1296) or 1295/1296.Therefore the odds are 1:1295


What are the chances of rolling double 6 3 times in a row?

The chance of rolling a double 6 is 1/6x1/6The chance of getting the 3 times in a row is(1/6x1/6)3 or (1/36)3 which is 1/46,656, a very small number!


What is the probability of rolling doubles three time in a row After rolling a dice 250 times?

the chances of rolling doubles once is 1 in 6; 3 times in a row it is 1 in 216. It does not make any difference after how many times you rolled the dice before.


What is th probability of rolling six three times in a row with a six sided dice?

Assuming that the die is a "normal" one (it has the numbers 1 to 6 and that it is fair), then the probability of rolling six three times in a row is 1/6*1/6*1/6 = 1/216 = 0.00463 The probability of rolling six three times in a row eventually is 1 (ie a certainty).


What is the probability of rolling 6 and four in backgammon three times in a row?

The answer depends on how many times in total the dice are rolled. As the total number of rolls increases, the probability rolling a 6 and 4 three times in a row increases towards 1.


What is the probability of rolling any number two times in a row with a die?

The probability of rolling any specific number on a six-sided die is ( \frac{1}{6} ). To find the probability of rolling that same number two times in a row, you multiply the individual probabilities: ( \frac{1}{6} \times \frac{1}{6} = \frac{1}{36} ). Thus, the probability of rolling the same number two times in a row is ( \frac{1}{36} ).


What are the odds of rolling doubles six times in a row?

Odds of rolling ONE six - 6:1 Odds of rolling TWO sixes - 36:1 Odds of rolling two sixes, SIX times - 216:1


What is the probability of rolling a five on one dime three times in a row?

When rolling a single die, the probability of rolling a five is ( \frac{1}{6} ). To find the probability of rolling a five three times in a row, you multiply the individual probabilities together: ( \left(\frac{1}{6}\right)^3 = \frac{1}{216} ). Therefore, the probability of rolling a five on one die three times in a row is ( \frac{1}{216} ) or approximately 0.00463.


What is the chance of rolling a 4 with a single die three times in a row?

The probability of rolling a 4 with a single six-sided die is 1/6. To find the chance of rolling a 4 three times in a row, you multiply the individual probabilities: (1/6) × (1/6) × (1/6) = 1/216. Therefore, the chance of rolling a 4 three times consecutively is 1 in 216, or approximately 0.46%.


Is it true that the probability of rolling a six with a fair die five times in a row is?

The probability of rolling a six with a standard die five times in a row is (1 in 6)5 which equals 1 in 7776 or about 0.0001286.

Trending Questions
Average weight for a 5' 10 man? How much silver in a 1922 silver dollar? What is the average hand span for a 11 year old? Sample letter for shifting unit to one place to another place? What is an average weight for a 13 year old girl that is 161 cm tall? What is the probability of rolling a pair of six-sided dice 25 times and getting double sixes once Give your answer as a decimal rounded to the nearest hundredth? What is a correct statement regarding materiality? A box contains two coins One coin is heads on both sides and the other is heads on one side and tails on the other One coin is selected from the box at random and the face of one side is observed If t? What best describes a bar graph? What is the level of measurement for family history of illness in statistics? What is one third plus two sixth? What is the probability of rolling a 4 5 times in a row on 1 die? How many sig figs are in 890 mL? Josh has 4 shirts white blue green and red and 4 pairs of socks in the same colors. Josh grabs a shirt and a pair of socks without looking. What is the probability that his shirt and socks will match? What is the most common definition of uneven? How many pounds is 1000 tons? How many people does it take to make a shirt? What is the percentage of men cheating on there girlfriends? Does it mean i'am pregnant if my period came twice in one month if it doesnt normally does that? What is the probability that a random selected poker hand contains exactly 3 aces given that it contains at least 2 aces?